### v1

# A Set-Sequence Model for Time Series

Abstract
In many financial prediction problems, the behavior of individual units (such as loans, bonds, or stocks) is influenced by observable unit-level factors and macroeconomic variables, as well as by latent cross-sectional effects. Traditional approaches attempt to capture these latent effects via handcrafted summary features. We propose a Set-Sequence model that eliminates the need for handcrafted features. The Set model first learns a shared cross-sectional summary at each period. The Sequence model then ingests the summary-augmented time series for each unit independently to predict its outcome. Both components are learned jointly over arbitrary sets sampled during training. Our approach harnesses the set nature of the cross-section and is computationally efficient, generating set summaries in linear time relative to the number of units. It is also flexible, allowing the use of existing sequence models and accommodating a variable number of units at inference. Empirical evaluations demonstrate that our Set-Sequence model significantly outperforms benchmarks on stock return prediction and mortgage behavior tasks. Code will be released.
1 Introduction
Important time series problems in finance and economics, such as estimating loan risk and forecasting stock returns, often involve predicting outcomes for a varying number of correlated units. For example, consider predicting the outcomes for a pool of 10,000 loans with 50 features per loan. A full joint model requires handling a total of 500,000 features to capture latent cross-sectional effects such as contagion [1, 2, 30]. This is practically intractable, so a common approach in the literature augments unit-level and macroeconomic feature series with manually crafted summary features intended to capture the latent cross-sectional effects (e.g. [22] employs lagged default rate for each zip code). Aside from requiring domain expertise, this approach does not scale well and does not reliably capture all relevant cross-sectional dynamics.
This paper proposes a Set-Sequence model for time-series, which learns the latent cross-sectional effects, eliminating the need to handcraft features. Figure 1 shows the architecture. Harnessing the exchangeability of units [^1] Exchangeability means the identity of individual units (e.g. stock ticker) is irrelevant – unit-level behavior is determined solely by observed features. Formally, for a problem of size $M$, the unit-level feature vectors $X^{i}$ are exchangeable: $P(X^{1},X^{2},\dots,X^{M})=P(X^{\pi(1)},X^{\pi(2)},\dots,X^{\pi(M)})$, for any permutation $\pi$. , the Set-Sequence model first computes an order-invariant set summary at each period, creating a time series of cross-sectional features. This series augments the time series for each unit separately, enabling the model to capture group dynamics without handcrafted features. Next, the prediction for each unit is made independently by a sequence model operating on this augmented time series containing both unit-level and cross-sectional features. This permits the use of any preferred sequence model (e.g. Transformer, RNN). Viewed through this lens, the Set-Sequence model makes simple architectural changes that have significant practical implications: it provides a way to extend a unit-specific model to include cross-sectional dependencies without needing to handcraft features, and the model allows inference over a variable number of units. Moreover, the compute scales linearly in the number of units. Section 3 provides details on our methodology.
Figure 1: The Set-Sequence model: The Set model estimates a cross-sectional summary at each time, in linear complexity over the number of units. This augments the original features of each unit and the Sequence model consumes this augmented series to make predictions for each unit independently. Note that the Set model has a look-back of $L$ time periods, where $L\geq 1$ is a model parameter. The number of units $M$ and time periods $T$ may vary at inference.
[…]
We next evaluate the Set-Sequence model on a stock portfolio prediction task using actual return data covering 36,600 US stocks. The model selects daily portfolio weights to maximize the portfolio’s annual Sharpe ratio (a widely used performance criterion). We compare against state-of-the-art sequence models and the leading task-specific CNN-Transformer of [11]. Over a 15-year out-of-sample period, the Set-Sequence model achieves between 22% and 42% higher Sharpe ratios compared to baseline models. These improvements are highly significant from an economic standpoint, and translate into substantial investment gains.
Finally, we evaluate the Set-Sequence model on a large-scale US mortgage risk prediction task, comparing it to a suite of sequence models and the leading task-specific deep learning model of [22]. The task involves classifying the monthly mortgage state (current, 30/60/90 delinquency, foreclosure, prepayment) given its history. The data cover 5 million loan-month observations, and contain 52 dynamic loan-level and macro-economic features. Efficiently harnessing cross-sectional patterns, the Set-Sequence model achieves significantly higher AUCs than the benchmark models. The model delivers more accurate risk metrics, enabling more effective risk management.
Our work underscores the importance of exploiting exchangeability in high-dimensional time series and shows that the Set-Sequence model is an effective tool for tackling this. Our model: (1) Outperforms strong benchmarks on both synthetic and real tasks. It accommodates a variable number of units at inference, which is of significant practical interest. (2) Efficiently addresses large problems due to its linear scaling along the cross-sectional dimension. (3) Is interpretable with a low-dimensional set summary offering a window into the dependency structure across units.
2 Related Literature
Traditional approaches to modeling multiple time series have relied on, for example, Vector Autoregressive (VAR) models [26], which scale quadratically in the cross-sectional dimension. Linear factor models [4, 29, 27, 28, 3] offer an improvement, greatly reducing dimensionality compared to classical VAR approaches. While these linear models struggle with non-linear dependencies, deep learning approaches to factor modeling [32] have emerged to address non-linearities, though they still require retraining when new units are introduced.
Recent deep learning architectures, while powerful, fail to exploit cross-sectional exchangeability while maintaining computational efficiency. Transformer variants [31, 41, 18, 40, 19, 42, 35, 20], State Space Models such as S4 [10], H3 [7], and others [39, 34], convolution based models (LongConv [8], Hyena [21], and Multilayer Perceptrons [6, 37] all scale quadratically in the cross-sectional dimension. This often necessitates modeling each unit individually with shared parameters, relying on handcrafted features to capture joint dependencies as in [22].
Some approaches have attempted to capture cross-sectional exchangeability more directly. [23] combine a local recurrent neural network (RNN) model with a global Gaussian Copula model, but their simple covariance structure may not capture non-linear joint dependencies and has only been demonstrated for single-feature units. Other work explores capturing global dependencies through low-rank matrix factorization trained iteratively alongside forecasting models [24], but this also needs refitting when new units are introduced. There is also research that exploits set structures, introduced by [38, 16], though it focuses on set representations across time rather than in the cross-section [12]. We provide a more in-depth literature review in Appendix A.
3 Set-Sequence Model
Consider a dataset $\mathcal{D}=\{({Y}^{i},{X}^{i})\}_{i=1}^{M}$ with $M$ units, where $Y^{i}=(Y_{1}^{i},Y_{2}^{i},\dots,Y_{T}^{i})\in\mathbb{R}^{d_{y}\times T}$ is the response and $X^{i}=(X_{1}^{i},X_{2}^{i},\dots,X_{T}^{i})\in\mathbb{R}^{d_{x}\times T}$ is a covariate. We allow data with different start and end dates but we assume the data are on a regular grid and padded so that all series have the same length. We use the notation $X^{i}_{(s,t)}=(X^{i}_{s},X^{i}_{s+1},\dots,X^{i}_{t})$ to denote the covariates from time $s$ to $t$ for unit $i$. Given a set of covariates $\{X_{(s,t)}^{j}\}_{j=1}^{M}$, the goal is to predict $Y^{i}_{t+1}$.
The Set-Sequence layer combines a set network for cross-sectional data processing with an arbitrary sequence-to-sequence layer for modeling temporal dependencies, as shown in Figure 1.
Set-Sequence Layer
Assume that $X^{1},\dots,X^{M}$ are the input units to the Set-Sequence layer, and $Y^{1},\dots,Y^{M}$ are the outputs of the layer. We use temporal chunks as many temporal patterns depend on several time steps to process. These inputs are processed in temporal chunks of size $L$, to allow the summary at the current period to use information from recent previous periods, producing feature embeddings via an embedding network, $\phi$. These embeddings are then fed through a set network $\rho$ to produce a permutation invariant, low dimensional, set summary $F_{t}$
$F_{t}=\rho\left(\frac{1}{M}\sum_{i=1}^{M}\phi\left(X_{(t-L,t)}^{i}\right)\right)\in\mathbb{R}^{r},$ (1)
We then concatenate the set summary to create an augmented feature $\tilde{X}^{i}$ that is fed individually into a given sequence layer SeqLayer
$\tilde{X}^{i}_{t}$ $=\psi([X^{i}_{t},F_{t}]^{T})$ (2)
${Y}^{i}_{t+1}$ $=\operatorname{SeqLayer}(\tilde{X}^{i}_{(1,t)}),$ (3)
for all $i\in\{1,\dots M\}$, and $t\in\{1,\dots,T\}$.
Set-Sequence Model
The Set-Sequence model replaces a subset of the layers in a sequence model with the Set-Sequence layer. For example, a Transformer, Long Convolution, SSM, or RNN layer can be used for $\operatorname{SeqLayer}$.
Alternative Layer Formulations
We use the linear set aggregation in Equation 1 due to its simplicity and empirically strong performance. As an alternative, we propose MHA-Seq, a multi-head attention layer that scales quadratically with the number of units. This layer is effective when the cross-sectional dimension is small or when cross-unit dependencies are particularly strong. For each unit, we compute a low dimensional set summary by using multi-head attention, by cross attending to all the other units, by the following equation
$F^{i}_{t}=\operatorname{MHA}\left([\phi(X^{1}_{(t-L,t)}),\dots,\phi(X^{M}_{(t-L,t)})]^{T}\right)_{i}.$ (4)
Other architectures are explored in Appendix B.
Implementation Details
Unless noted, we use LongConv [8] for the sequence layer, two-layer FFNs with dropout for $\phi$, $\rho$, and $\psi$, and chunk size L = 3. MHA-Seq differs only by using five attention heads. Full hyperparameters are in Appendix C.
4 Baselines and Task Selection
Task Selection
Common multivariate-forecasting benchmarks—ETTh1, ETTm1, Exchange, and Traffic (see details in [41])—are ill-suited for our study: their unit counts are small enough that a full joint model remains tractable or they lack a clear exchangeable-unit structure. We therefore evaluate on three datasets—a synthetic contagion setting, an equity-portfolio regression task, and a mortgage-risk classification task—each defined over exchangeable units. They contain 1000, 500, and 2500 units with 4, 79, and 52 features per unit, giving cross-sectional dimensions of 4000, 39500, and 130000 that meaningfully stress models built for high-dimensional cross-sections.
General and Domain-Specific Baselines
We benchmark against five widely adopted sequence architectures— Transformer [31], S4 [9], H3 [7], Hyena [21], and LongConv [8]. Together they span attention-, state-space-, and convolution-based methodologies, each of which our Set-Sequence layer can augment. We treat them as task-agnostic baselines on all three datasets. Because these models share a common backbone structure, we hold input/output dimensions, hidden size, learning-rate schedule, and other hyperparameters fixed, so the only variation lies in the sequence layer (and whether a Set module is present). This yields a fair comparison across models. Each task section also reports results for the strongest task-specific domain baseline. Hyperparameters for all models are in Appendix C.
5 Synthetic Task
This section examines the performance of the Set-Sequence model in a setting similar to default risk prediction for corporate and other loans, where contagion has been shown to play an important role [1, 2, 30]. The goal is to predict the transitions probabilities in the next time step for each of the units.
Task Description
We simulate 1000 exchangeable units, each with binary feature $x\in\{0,1\}$ over 100 time steps. Units transition between three states (state 3 is absorbing/default) with transition matrix proportional to:
$\begin{pmatrix}1+x&1&(\lambda_{x,t}+\mu)(1+0.1x)\\ 1&1+x&(\lambda_{x,t}+\mu)(1+0.1x)\\ 0&0&1\end{pmatrix}$ (5)
Default intensity evolves as $\lambda_{x,t+1}=\beta\lambda_{x,t}+\alpha N_{x,t}$, where $N_{x,t}$ is the fraction of type-$x$ units defaulting at time $t$. Parameters $\mu=0.001$, $\alpha=4$, and $\beta=0.5$ yield an average default rate of $\sim$1%. The model captures contagion via $\lambda_{t}=(\lambda_{0,t},\lambda_{1,t})$, where defaults increase future default risk. The core challenge in this synthetic setting is modeling joint dependencies among many units via the low-dimensional latent factor $\lambda_{t}$, itself a function of each unit’s state.
5.1 Comparison with Sequence Models
[…]
Table 1: Comparison of Set-Seq model with Sequence Models operating on one unit at a time (single) as well as the full cross-section (joint).
[…]
LongConv (Joint) 0.0369 0.681
S4 (Joint) 0.0375 0.676
H3 (Joint) 0.0401 0.675
Transformer (Joint) 0.0357 0.506
Hyena (Joint) 0.0358 0.702
LongConv (Single) 0.00182 0.757
S4 (Single) 0.00155 0.758
H3 (Single) 0.00170 0.751
Transformer (Single) 0.00156 0.758
Hyena (Single) 0.00169 0.760
Set-Seq (Ours) 0.00028 0.801
Model Ablations
We conduct an ablation study of the Set-Sequence model on the synthetic task, with results summarized in Table 2. We vary the cross-sectional modeling component (None, Set, or Multi-Head Attention) and the sequence length (1 or 50) to assess their impact on predictive accuracy and computational cost. The largest performance gain comes from adding a set model for the cross-sectional dimension, compared to modeling units independently. While replacing the set model with Multi-Head Attention (MHA) further improves accuracy, it incurs a $\times3$ increase in training time and a 3.5×3.5$\times$3.5 × increase in memory usage, due to the quadratic scaling of attention with the number of units.
Table 2: Ablation study of the Set-Sequence model on the synthetic task, varying the cross-sectional modeling (None, Set, or Multi-Head Attention) and sequence length (1 or 50). Sequence Length refers to the model’s effective lookback or kernel length. Relative Epoch Time and Relative Max Train Memory are reported relative to the Set-Sequence model with Set and Sequence Length 50. All models were trained on an Nvidia RTX A6000 GPU.
Seq. Model Set Model Seq. Len. KL(true|predicted) AUC Rel. Epoch Time Rel. Mem
LongConv None 1 0.0021 0.693 0.51 0.86
LongConv None 50 0.0018 0.757 0.64 0.93
LongConv Set 1 0.00054 0.792 0.60 0.88
LongConv Set 50 0.00028 0.801 1.00 1.00
LongConv MHA 50 0.000044 0.815 3.35 3.52
5.2 Generalization Across Number of Input Units
Task
Given the setting with $M$ units as described above, the task is to predict the next state transitions accurately based on the previously seen data. We observe $\bar{N}_{t}$ which is the fraction of units defaulted when only $n$ out of $M$ units are observed, hence a noisy realization of $N_{t}$. In the default prediction setting, this correspond to having a large universe of loans evolving jointly, but only observing a limited subset of the data.
Baseline Based on the Kalman Filter
We develop a strong baseline that represents near ideal performance on the task when trained with infinite data. We assume that the true transition matrix is known, and we also assume that the dynamics of $\lambda_{t}$, $\lambda_{t+1}=\beta\lambda_{t}+\alpha N_{t}$ are known, with the only unknown factor being not observing the true fraction of transition to the absorbing state $N_{t}$, but we only observe $\bar{N}_{t}$, where $\bar{N}_{t}$ is the observed fraction given n out of M units. The problem is now to infer the best estimate of $\lambda_{t}$ based on partial observations $\bar{N}_{1},\dots,\bar{N}_{t}$. In Appendix D.3 we derive an estimation $\hat{\lambda}_{t}$ that is near optimal, based on a Kalman Filter. This estimate is then used along with the true transition dynamics to generate transition probabilities for the next time period for each unit. The goal is hence to achieve performance as close as possible to the Kalman Filter baseline.
Model Training
During training we draw the number of units per batch as
$D\sim(1-\gamma)\cdot\delta_{M}+\gamma\cdot\text{LogUniform}(1,M),$
with $\gamma=0.08$; Thus most batches use all $M=1000$ units while a few expose smaller pools for generalization. An analysis of other sampling schemes along with more training details are provided in Appendix D.
Metrics
We evaluate models on their ability to predict transitions to the absorbing (rare) state using three metrics: AUC, correlation, and $R^{2}$. The AUC measures the model’s ability to rank true absorbing transitions above non-absorbing ones. Correlation captures the linear agreement between predicted and true transition probabilities. $R^{2}$ quantifies the variance explained for the absorbing transitions specifically. See Appendix 5.2 for formal definitions.
Results
Figure 2: Comparison between the Set-Sequence model, MHA-Sequence model, and the Kalman Filter, by considering the AUC, $R^{2}$, and Correlation, each for the absorbing (rare) state. The Set-Sequence model reaches near the oracle-like Kalman Filter model across the full range of units for inference, for all the considered metrics.
In Figure 2, the performance of both the Set-Sequence and the MHA-Seq model are compared with the oracle-like Kalman Filter baseline for inference sizes ranging from one unit (out of 1000) to the full information setting with 1000 units. We note that both the Set-Sequence and the MHA-Seq models reach near optimal performance on the task for all considered metric, across the full spectrum of number of input units. This shows that a single model generalizes at inference to unseen or rarely seen input sizes from 1 to 1000 units. An important consideration is that the oracle-like Kalman Filter baseline are computed using the infinite data limit, and the set sequence model is trained on a restricted size training set. The MHA-Seq offers a slight performance advantage on the synthetic task, but this comes at a computational disadvantage due to quadratic scaling in the number of input units.
Interpretability
The Set-Sequence model produces interpretable results. Table 3 shows the average correlation across 100 samples of length 100 from the test set, between the first set variable in layer i of the Set-Sequence and the latent factor $\lambda_{0,t}$. As we observe more units the correlation increases, with a very high average correlation $0.951$ in the fully observed case in layer 3. Appendix D.2 contains visualizations of the learned set variables.
Table 3: Average correlation for 100 samples, each with 100 time steps between the set summary at layer i and the true latent variable, $\lambda_{0,t}$. Best values in each column are in bold, with the second best underlined.
# obs 20 50 100 200 500 1000
Layer 1 0.234 0.237 0.236 0.233 0.233 0.232
Layer 2 0.160 0.165 0.165 0.167 0.165 0.166
Layer 3 0.349 0.478 0.576 0.699 0.860 0.951
Layer 4 0.263 0.330 0.414 0.547 0.642 0.690
Layer 5 0.444 0.520 0.610 0.681 0.759 0.804
6 Empirical Application: Equity Portfolio Prediction
We evaluate the Set-Sequence model on a portfolio prediction task. In the task, given a universe of $N$ assets, at the current time, the model outputs $N$ portfolio weights, indicating the portfolio we invest in. The portfolio returns are realized from the inner product between the next day returns and the chosen portfolio weights. The goal is to optimize the Sharpe ratio, a measure of the risk-adjusted returns. This is a common objective in the asset pricing literature, as in [11].
Dataset
We use data from the Center for Research in Security Prices (CRSP). The full dataset consists of 36600 stocks. For each period, we train on 8 years and test the model on the following year. For each model fitting, we filter out stocks where any returns in the train or test period are not available. After this filter, we filter to include only the 500 assets with the largest market cap at the last quarter of the train period, mimicking the S&P 500 index. This ensures that we consider assets that are liquid with small bid-ask spreads, aligned with prior work [11]. We use a subset of the features from [5], with the same normalization (cross-sectional rank normalization). The features are based on recent cumulated returns, volatility, volume, and quarterly firm characteristics (such as book-market ratio, and total assets). In total we use 79 features listed in Table 10 and Table 11 in the Appendix.
Objective
Let $X^{i}_{t}$ be the covariates at time t, and let $Y^{i}_{t}$ be the excess daily returns for asset i (returns with the risk free rate subtracted), where $X_{t}$ is the covariates for all assets, and $Y_{t}$ is the returns for all assets. We then use $X_{1},\dots,X_{t-1}$ to predict $\hat{w}_{t}$, the portfolio weights at the next step. Given a series of daily excess returns, $r_{1},\dots,r_{T}$, where $r_{t}=Y_{t}^{T}\hat{w}_{t}$, the daily Sharpe ratio (SR) is then
$SR=\frac{\bar{r}}{\sqrt{\frac{1}{T}\sum_{t=1}^{T}(r_{t}-\bar{r})^{2}}},$ (6)
where $\bar{r}=\frac{1}{T}\sum_{t=1}^{T}r_{t}$. We $L_{1}$ normalize the portfolio weights at each timestep to maintain unit leverage ($||w_{t}||_{1}=1$). Aligned with [11] we use the negative of the Sharpe ratio as the optimization objective.
Baselines
We benchmark against the sequence baselines described in Section 4. Each baseline is trained on a single unit at a time with weights shared across units—a strategy that generalizes far better than feeding the entire cross-section, which performs poorly (see Section 5.1). We further compare with task-specific models: the CNN-Transformer and the additional baselines of [11], which use aligned dataset, time span, and feature set (the latter following [5]).
Metrics
The main metric to evaluate the performance is the annualized Sharpe ratio. We also give the percentage yearly return, the yearly return volatility, the average daily turnover, defined as the average $L_{1}$ distance between the weights in consecutive days. Beta is the covariance of portfolio and market returns divided by the market variance, measuring market exposure. A low beta is desirable to ensure returns are independent of market direction.
Results
Table 4 shows the aggregate performance. The Set-Sequence model outperforms every sequence model baseline. We note a high annualized mean Sharpe ratio over the full period of 4.82, 22% higher than the second best sequence model (S4), and 32% higher than the LongConv model, which the Set-Sequence model use for the sequence component in the experiment. Compared with the other sequence models, the Set-Sequence model shows robust results, with a Sharpe ratio standard deviation over 5 random initialized training runs of 0.12, compared with 0.52 for the Transformer and 0.29 for S4.
Figure 3 shows the cumulative returns for the models. We also compare the Set-Sequence model with domain specific models on the time period Jan. 2002 - Dec. 2016 to be aligned with time period of prior work. Table 5 reports that the general Set-Sequence model outperforms the CNN-Transformer designed specifically for the equities task by 42 % in terms of the Sharpe ratio. In Section E in the Appendix we include an analysis of the Sharpe ratio when we include transaction costs. We use a net Sharpe objective function and show that the Set-Sequence model still outperforms all baselines in the presence of transaction costs.
[…]
Model Sharpe Return % Std Dev % Beta Daily Short
Ratio Return Turnover Fraction
[…]
Set-Seq (Ours) $\textbf{4.82}\,\pm\,0.12$ 13.0 2.69 0.028 0.91 0.48
Table 5: Out-of-sample annualized performance metrics from 2002 to 2016. All CNN-Transformer models are based on the IPCA method described in [11]. For the Set-Sequence model the results are the mean over 5 training initialization seeds. For the Sharpe we also report its standard deviation over the seeds.
Model Sharpe $\mu$ [%] $\sigma$ [%]
CNN+Trans K=5 4.16 8.7 2.1
CNN+Trans K=8 3.95 8.2 2.1
CNN+Trans K=10 3.97 8.0 2.0
CNN+Trans K=15 4.17 8.4 2.0
Fourier+FNN K=10 1.93 7.6 3.9
Fourier+FNN K=15 2.06 7.9 3.8
OU+Thresh K=10 0.86 3.1 3.6
OU+Thresh K=15 0.93 3.2 3.5
[…]
Figure 3: Cumulative returns from 2002 to 2021 for the Set-Sequence model and sequence model baselines. We also show the equally weighted market portfolio of the largest 500 stocks by market cap.
7 Empirical Application: Mortgage Risk Prediction
Next we evaluate the Set-Sequence model on a mortgage risk prediction task, with the goal to predict the state (current, 30 days delinquent, 60 day delinquent, 90+ days delinquent, Foreclosure, Paid Off, and Real Estate Owned) in the next month given a history of 52 covariates (FICO score, loan balance, current interest rate etc.) for a mortgage.
Dataset
We source the data from 4 ZIP codes in the greater LA area as a subset of the CoreLogic Loan-Level Market Analytics Dataset, as these are among the ZIP codes with most active mortgages. We follow the same data filtering procedure as in [22], resulting in 5 million loan month transitions from 117523 active loans. We use the following train/validation/test split: Jan. 1994 - Jun. 2009/Jul. 2009 - Dec. 2009/Jan. 2010 - Dec. 2022. The number of active loans in the dataset varies between 4000 in Jan 1994, up to 22000 in 2006. In Appendix F.1, we include additional dataset descriptions, filtering procedure, transition probabilities and counts of the number of active loans over time.
Figure 4: AUC gain (Set-Seq – baseline [22]) for each transition. Deeper green indicates a larger Set-Seq advantage; red indicates the baseline performed better. Larger circles indicate more common transitions, with size proportional to the log of their frequency. Only transitions occurring at least 10 times are included.
Figure 5: Foreclosure rate over the dataset, as well as the learned set representation in the first set layer in the neural network. The Set-Sequence model learns to predict the foreclosure rate for joint default modeling.
Setup
We compare the Set-Sequence model with the domain specific models presented in [22], representing the current state of the art, along with the sequence models from Section 4. The models from [22] includes a neural network baseline has 5 layers with early stopping and dropout regularization, and a logistic regression baseline. Both baselines use the features at the current time to predict the state in the following month. For each training batch, for a given time we sample 2500 loans out of all active ones. We use a sequence length of 50, and use a multi-class cross entropy loss during training.
Results
Table 6: Comparing performance of different models on the mortgage risk prediction task on the test set. The average AUC is the mean of the AUCs to go between different states. The average transition probabilities are the average probability for the correct class over the full test set.
Model Cross Entropy Avg. AUC
5-Layer NN 0.205 0.642
Log. regression 0.225 0.622
LongConv 0.216 0.669
S4 0.226 0.681
H3 0.222 0.583
Transformer 0.227 0.666
Hyena 0.213 0.674
Set-Sequence (Ours) $\mathbf{0.200}$ $\mathbf{0.683}$
In Figure 5 we see that the Set-Sequence model robustly outperforms the prior state of the art model on the task, with better AUC for 22 out of 25 transitions, including the most common ones and economically important transitions, such as those to the paid-off state from current. In Figure 5, the foreclosure rate, a known source of cross unit dependency, is shown to have a Pearson correlation of 0.67 with the set summary in the first Set-Sequence layer, indicating the interpretability of the learned set summaries. Table 6 shows that the Set-Sequence model achieves the best performance on both cross-entropy loss and average AUC, outperforming all domain-specific and general sequence model baselines. In particular, it improves average AUC by 4 points over the best domain-specific baseline (5-layer neural network).
8 Conclusion
This paper introduced the Set-Sequence model, an efficient architecture for capturing latent cross-sectional dependencies without manual feature engineering. By learning a shared cross-sectional summary at each period with a Set model and subsequently augmenting individual unit time series for a Sequence model, our approach demonstrates strong empirical performance on synthetic, stock return, and mortgage risk tasks, significantly outperforming benchmarks. While many existing multivariate models are powerful, they are often not natively designed to exploit unit exchangeability in multivariate financial time series. The Set-Sequence model offers a flexible and scalable solution.
Limitations: Future work should include a deeper exploration of alternative set architectures within the Set model, an investigation into the model’s performance on datasets with weaker exchangeability assumptions, and further analysis of the interpretability of the learned set summaries.
Appendix A Additional Literature Review
Many problems exhibit cross-sectional exchangeability: the identity of each unit (e.g., stock, loan) is irrelevant—behavior depends only on observed features. Traditional multivariate time-series models can still be applied in two ways: (i) individual modeling, where one unit is processed at a time with shared parameters, or (ii) a full joint model, which is oftentimes infeasible.
Individual Unit Modeling
This approach naturally handles a variable number of units and allows immediate inference for unseen units because the same parameters are shared, but it ignores cross-unit dependencies. Representative methods include classical Vector Autoregressive (VAR) models [26], which scale quadratically in the cross-section, and a broad class of modern deep-learning time-series models: Transformer variants [41, 18, 42, 35, 20], state-space models [10, 39, 34], convolutional models [8], multilayer perceptrons [6, 37], and graph neural networks [36, 25].
Exchangeability with Fixed Input Size
These models treat units as exchangeable but must be retrained when new units appear.
  1. 1.
    Dynamic factor models. Linear versions [4, 29, 27, 28, 3] reduce dimensionality but require re-estimating loadings (and sometimes factors) whenever units change. Deep factor models [33] replace linear dynamics with RNNs, yet still demand refitting loadings and use local linear/Gaussian components that may miss nonlinear effects.
  2. 2.
[…]
Exchangeability with Dynamic Number of Units
[…]
Financial Applications
Joint modelling is vital in finance—for instance, contagion in loan defaults [1, 2, 30]. Previous work often operated on one unit at a time [22, 14]. In portfolio optimisation, models typically impose a fixed ordering and limit cross-sectional width [15], overlooking exchangeability.
Appendix B Set-Seq model with Gated Selection
This section describes an alternative modeling approach we found to be promising. The Gated Selection cross-sectional modeling approach is computationally in between the cost for the full attention and the linear set aggregation in the cross section.
Gated Selection Layer
We use temporal chunks as many temporal patterns depend on several time steps to process. These inputs are processed in temporal chunks of size $L$, to allow the summary at the current period to use information from recent previous periods, producing feature embeddings via an embedding network, $\phi$. Then, we define a gating matrix G by the following equations
$A_{ij}=\frac{X_{i}^{T}W^{T}WX_{j}}{\|WX_{i}\|_{2}\|WX_{j}\|_{2}}$
$G=\text{Softmax}(A),$
where the Softmax is taken along each row. Finally, we produce a per-unit, low-dimensional set summary $F_{t}^{i}$ by
$F_{t}^{i}\;=\;\rho\!\Bigl{(}\,\sum_{j=1}^{M}\,G_{i,j}\,\phi\bigl{(}X_{(t-L,\,t)}^{j}\bigr{)}\Bigr{)}\;\;\in\;\mathbb{R}^{r},$
where $\rho(\cdot)$ is an output projection. The gating, with $G_{i,j}\in[0,1]$ allows the summary statistic of unit i to mainly depend on the other units with ’similar’ features to unit i. We recover the standard Set-Seq model when we set $G_{i,j}=\frac{1}{M}$, for all $i,j$ where $M$ is the number of units. We provide experiments using the Gated Selection mechanism in Section 14.
Similarity with Attention
The Selection mechanism shares important similarities with attention, but is different in the following ways:
  • It takes the key and query matrix to be the same (which makes sense given the unit exchangability)
  • It replaces the $\sqrt{d}$ normalization with a norms of $WX_{i}$ and $WX_{j}$, also motivated by unit exchangeability, where all should have similar norms.
  • Most importantly, we only compute one G for all times for the sample, this has a large impact of the memory usage, removing the linear scaling with sequence length needed for gradient computations in the selection mechanism.
Appendix C Experiment Details
We here detail the hyperparameters used for training the Set-Sequence model. The hyperparameters for the synthetic task, mortgage risk task, and equity portfolio prediction task are shown in Table 7. In Table 8 we show the layer hyperparameters for the sequence model baselines used. The number of layers, learning rate schedule, learning rate, and other training parameters are the same as for the Set-Sequence model. The hyperparameters used are aligned with baseline hyperparameters from the corresponding papers.
Table 7: Hyper-parameter settings for each task.
Hyper-parameter Synthetic Mortgage Equity
Set-Seq layers 5 6 6
Sequence layers 1 0 0
$d_{\text{model}}$ 800 300 64
Learning rate 0.003 0.003 0.003
Dropout 0 0.10 0.10
Epochs 40 40 30
# Samples 250
Time steps 100 50 246
Number of units 1–1000 2500 500
Sequence model LongConv LongConv LongConv
Chunk size $L$ 3 3 3
Set summary dim 2 2 2
$\phi$ output dim 5 5 5
Conv. weight decay 0 0.05 0.05
Table 8: Hyperparameter settings for the five sequence model baselines. For all models we use six layers, and the hidden dimension for a given task is the same as for the Set-Sequence model. The learning rate schedule, optimizer, and other training parameters are the same as for the Set-Sequence model.
LongConv S4 H3
Kernel Length 30 $d_{\text{state}}$ 64 $d_{\text{state}}$ 64
Dropout 0.0 Learn $\theta$, $a$ True Head Dim 1
Nr Layers w/ Set 6 Skip Connection True Mode diag
Set Embed Dim 5 Learning Rate 0.001 Measure diag-lin
Learning Rate 0.001
Hyena Transformer
Order 2 Heads 8
Filter Order 64 Causal True
Num Heads 1 Learning Rate 0.001
Dropout 0.0 Dropout 0.0
Compute resources
All the experiments in the paper was conducted on a Linux cluster with 5 NVIDIA RTX A6000 GPUs, each with 49140 MB memory, running on CUDA Version 12.5. The cluster has 256 AMD EPYC 7763 64-Core Processor CPUs. For the synthetic task, each model was trained for less than 1 hour on 1 GPU. For the equities task, each train run (one run per validation year and random seed) took around 30 minutes on a single GPU. The experiments for the mortgage risk prediction task ran for around 1 hour per model training on one GPU.
Data Sources
For the mortgage risk case study we use the Version 2.0 CoreLogic® Loan-Level Market Analytics dataset released on July 5, 2022. For the equities portfolio prediction task we use base data derived from Center for Research in Security Prices, LLC (CRSP).
Appendix D Synthetic Task: Additional Details and Results
D.1 Ablations on the training input units sampling
As an ablation we consider sampling the number of units per batch with $D\sim\delta_{k}$, i.e. the number of input units is always k. The results are shown in Figure 6, where we use, 10, 100, 1000 samples (out of a total of 1000 units), compute adjusted, meaning that we train for 100 times more epochs (as each sample has 100 times fewer units) in the $n=10$ case compared with $n=1000$ observed units. We see a significant deterioration in performance when we only sample with 10 units, indicating that this is not sufficient to learn the joint structure in the data. The trade-off is that we get slightly better performance when we do inference on the same number of units that we trained on, but the generalization across units is significantly worse. We see that these methods provide worse generalization performance compared with the $(1-\gamma)\delta_{1000}+\gamma\operatorname{Logunif}$ sampling. We choose the LogUniform distribution as this gives roughly the same amount of samples in $(2,4),(4,8),\dots(500,1000)$, each can be viewed as a different data richness regime.
Figure 6: The performance for the MHA-Seq model depending on the training sampling method. $\delta_{k}$ denotes that all samples have input k units. The method are compute adjusted so that the number of epochs are scaled up to 400 (vs 40) for the case with input size 100, and scaled up to 4000 for input size 10.
D.2 Interpretability
In Figure 7 we show an example of the set summary for several different number of observed units for one sample in the test set. As expected, the correlation increases as we observe more units, with the sample having $98\%$ correlation between the set summary in layer five with the true joint effect $\lambda_{0,t}$. Also note how the set summary in layer 5 learns a more smooth representation compared with layer 3.
Figure 7: On the synthetic task with 1000 units in total, the Set-Sequence model’s set summary in layer 3 and layer 5, learns the true joint effect $\lambda_{0,t}$, showing the interpretability of the Set-Sequence model. We see that as the number of observed units approach the fully observed case (1000 units), the set summaries become more aligned with the true effect, with a high correlation ($r=0.98$) between the true joint effect and the set summary in layer five.
D.3 Kalman Filter Baseline
The synthetic task is described in Section 5. We assume that the transition matrix is known, and we also assume that the dynamics of $\lambda_{t}$, $\lambda_{t+1}=\beta\lambda_{t}+\alpha N_{t}$ are known, with the only unknown factor being not observing the true fraction of rare events $N_{t}$, but we only observe $\bar{N}_{t}$, where $\bar{N}_{t}$ is the observed fraction of rare events, only having access to n out of M units. The problem is now to infer the best estimate of $\lambda_{t}$ based on a partial observation of the fraction of rare events. We utilize a Kalman Filter [13] to estimate $\lambda_{t}$ with $\hat{\lambda}_{t}$. The resulting filter estimates
$\hat{\lambda}_{t+1}=\beta\hat{\lambda}_{t}+\alpha E[N_{t}](1-K(t))+\alpha K(t)\bar{N}_{t},$
a convex combination of the expected fraction of transitions to the absorbing state and the observed fraction of transitions to the absorbing state, where the Kalman Gain $K(t)$, is reflects how much noisily $\bar{N}_{t}$ estimates $N_{t}$. $K(t)=1$ indicates that we fully observe all units, whereas $K(t)=0$ correspond to the setting where none of the units are observed. The Kalman based baseline then uses $\hat{\lambda}_{t}$ to compute the estimated transition matrix 5, giving the predicted distribution of states in the next period. The Kalman Filter baseline assumes the true dynamics are known, whereas our model needs to learn the dynamics from a finite data. This provides an approximate upper bound for data-driven models.
Approximate Dynamical Model
We utilize that $N_{t}$ is a sum of $M$ Bernoulli random variables, and can hence be approximated with a normal distribution, where
$N_{t}$ $\sim\mathcal{N}(p_{t},\sigma_{N_{t}}^{2}),$
$\sigma_{N_{t}}^{2}$ $=\frac{p_{t}(1-p_{t})}{M},$
$p_{t}$ $=\frac{\lambda_{t}+\mu}{2+x+(\lambda_{t}+\mu)(1+0.1x)}(1+0.1x),$
Here, $p_{t}$ is the rare event transition probability, which is the expected fraction of rare events. We don’t observe $N_{t}$ but rather $\bar{N}_{t}$, which approximately follow, by the Central Limit Theorem:
$\bar{N}_{t}=N_{t}+\epsilon_{t},$
where
$\epsilon_{t}$ $\sim\mathcal{N}(0,\sigma_{\epsilon_{t}}^{2}),$
$\sigma_{\epsilon_{t}}^{2}$ $=\alpha^{2}p_{t}(1-p_{t})\left(\frac{M-n}{M^{2}}+\frac{(M-n)^{2}}{M^{2}(n+1)}\right).$
Kalman Filter
We aim to estimate the parameter $\lambda_{t+1}$ using a Kalman Filter, where there are $M$ units in total and $n$ are observed. The high-level intuition is that to update $\lambda$, we want to use our current estimate of $\lambda$ and a convex combination of the expected value for $N_{t}$ and the partially observed value of $N_{t}$, where we put more weight on the observed value the higher n, the number of observed units, takes. Let $P_{t}$ be the variance of the estimate for $\lambda_{t}$. The Kalman update equations are then given by the prediction step
$\hat{\lambda}_{t+1|t}$ $=\beta\hat{\lambda}_{t}+p_{t},$
$P_{t+1|t}$ $=\beta^{2}P_{t}+\sigma_{N_{t}}^{2},$
Kalman Gain computation
$K_{t}=\frac{P_{t+1|t}}{P_{t+1|t}+\sigma_{\epsilon_{t}}^{2}},$
and the update step
$\hat{\lambda}_{t+1}$ $=\hat{\lambda}_{t+1|t}+K_{t}\left(\bar{N}_{t}-p_{t}\right),$
[…]
$P_{t+1}$ $=(1-K_{t})P_{t+1|t}.$
These equations provide a way to incorporate information from both the expected fraction of transitions to the absorbing state at time t, as well as the partially observed fraction of transitions to the absorbing state at time t.
Visualizing the Kalman Filter
Figures 8 show the Kalman Filter estimates of $\lambda_{x,t}$ for $x\in\{0,1\}$, alongside the estimated gain $K(t)$ and the state estimate variance $P(t)$. In both cases, $n=500$ corresponds to full observation (since 500 units have each feature value), resulting in zero estimation variance and $K(t)=1$, as expected. As $n$ decreases, estimation variance increases and the Kalman gain decreases, causing the estimate to rely more on the prior mean than the noisy observations. This degrades tracking accuracy, as seen in the right panels. The transition model uses $\mu=0.001$, $\alpha=4$, and $\beta=0.5$.
Figure 8: Kalman Filter estimates of $\lambda_{x,t}$, state estimate variance, and Kalman gain $K(t)$ for $x=1$ (top) and $x=0$ (bottom). Both cases use 500 total units per feature value. As $n$ decreases, estimation variance increases and the Kalman gain $K(t)$ drops, causing less accurate tracking of $\lambda_{x,t}$. The transition model uses $\mu=0.001$, $\alpha=4$, and $\beta=0.5$.
Improvement of Kalman Filter over simple estimate
Here we compare the Kalman Filter with other simple baseline models. Figure 9 shows the performance of the Kalman vs the baseline that sets a fixed gain $K(t)=1$, meaning that it estimates $\lambda_{t}$ with $\tilde{\lambda}_{t}$, where
$\tilde{\lambda}_{t+1}=\beta\tilde{\lambda}_{t}+\alpha\bar{N}_{t}$, i.e. it uses the observed fraction of transitions to the absorbing state to estimate $\lambda_{t}$ and does not consider the expected number of default events.
Figure 9: The Kalman Filter vs "Partially Observed Hawkes" which is the Kalman model where $K(t)$ is fixed equal to 1.
Appendix E Equity Portfolio Prediction: Additional Details and Results
Table 9: Out of sample annualized performance metrics net transaction costs 2002 - 2016, trained with a net sharpe ratio objective. Details of the models other than the Set-Sequence model are found in [11]. The training setting for the sequence models and Set-Sequence model is in Appendix C.
Model Net Sharpe Return % Std Dev %
CNN+Trans K=0 (IPCA) $0.52$ $8.5$ $16.3$
CNN+Trans K=1 (IPCA) $0.85$ $5.9$ $6.9$
CNN+Trans K=3 (IPCA) 1.24 $6.6$ $5.4$
CNN+Trans K=5 (IPCA) $1.11$ $5.5$ $5.0$
CNN+Trans K=10 (IPCA) $0.98$ $5.1$ $5.2$
CNN+Trans K=15 (IPCA) 0.94 $4.8$ $5.1$
LongConv $1.64$ $7.41$ $4.52$
S4 2.17 $8.31$ $3.82$
H3 $0.76$ $3.71$ $4.89$
Transformer $1.59$ $7.40$ $4.66$
Hyena $0.60$ $3.59$ $6.00$
Set-Seq Model (ours) 2.46 $8.4$ $3.4$
Transaction Cost Analysis
We test the Set-Sequence model and the other baseline models under realistic transaction costs. In particular, for each day we subtract the transaction costs from the daily portfolio returns, which are used to compute the Sharpe ratio. Following [11], we use the following model for the transaction cost for portfolio weights $w_{t}$ at time t, and $w_{t-1}$ at t-1:
$\text{cost}(w_{t},w_{t-1})=0.0005\times||w_{t}-w_{t-1}||_{1}+0.0001\times||max(-w_{t},0)||_{1}.$
The return at time t, can then be expressed as $r_{net,t}=r_{t}-\text{cost}(w_{t},w_{t-1})$. We now train the Set-Seq model and the Sequence model baselines with a modified objective using the net return Sharpe as the loss function. In Table 9 we show the net Sharpe ratio after costs in Jan. 2002 - Dec. 2016. We see the gain from modeling the cross-section as the set-sequence model outperforms all the sequence models in terms of net Sharpe Ratio, as well as the domain specific CNN-Transformer.
Equity Task Features
Table 10 shows the rank normalized features. In addition to a subset of the features in [5] we include features for daily return, weekly return, weekly volatility, daily volume. In order to keep the level information, we also include cross-sectional median features in Table 11, following [5].
Table 10: Firm-specific characteristics (six categories) used as features in the equity-portfolio-optimization task. Construction details are in the Internet Appendix of [5].
Past Returns Value
r2_1 Short-term momentum A2ME Assets to market cap
r12_2 Momentum BEME Book-to-market ratio
r12_7 Intermediate momentum C Cash + short-term inv. / assets
r36_13 Long-term momentum CF Free cash-flow / book value
ST_Rev Short-term reversal CF2P Cash-flow / price
Ret_D1 Daily return Q Tobin’s Q
Ret_W1 Weekly return Lev Leverage
STD_W1 Weekly volatility E2P Earnings/Price
Investment Trading Frictions
Investment Investment AT Total assets
NOA Net operating assets LME Size
DPI2A Change in PP&E LTurnover Turnover
Rel2High Closeness to 52-week high
Resid_Var Residual variance
Profitability Trading Frictions (cont.)
PROF Profitability Spread Bid–ask spread
CTO Capital turnover SUV Standard unexplained volume
FC2Y Fixed costs / sales Variance Variance
OP Operating profitability Vol Weekly Trading volume
PM Profit margin RF Risk-free rate
RNA Ret. on net operating assets Beta Beta with market
D2A Capital intensity
Intangibles
OA Operating accruals
OL Operating leverage
PCM Price-to-cost margin
Table 11: Cross-sectional median firm-characteristic variables and their stationary transformations ($t\text{Code}$).  Each feature is the cross-sectional median of the underlying firm characteristic. Note: The transformations ($t\text{Code}$) are (1) no transformation; (2) $\Delta x_{t}$; (3) $\Delta\log(x_{t})$; (4) $\Delta^{2}\log(x_{t})$.
Variable $t$Code Variable $t$Code Variable $t$Code Variable $t$Code
A2ME 3 AT 4 BEME 3 Beta 1
C 3 CF 2 CF2P 3 CTO 3
D2A 3 DP2A 3 E2P 3 FC2Y 3
Investment 3 Lev 3 LME 4 LTurnover 3
NOA 3 OA 2 OL 3 OP 3
PCM 3 PM 3 PROF 3 Q 3
Ret_D1 2 Rel2High 3 Resid_Var 3 RNA 3
r2_1 2 r12_2 2 r12_7 2 r36_13 2
Spread 3 ST_REV 2 SUV 1 Variance 3
Vol 4 STD_W1 3 Ret_W1 2
Appendix F Mortgage Risk Prediction: Additional Details and Results
F.1 Additional Data Descriptions
We provide additional details about the CoreLogic Loan-Level Market Analytics Dataset. The dataset is over 870 GB and contains month-by-month transition data for mortgages in the US. The dataset starts in 1988, and has data to 2024, with data from 30000 ZIP codes in the US. We restrict ourselves to the top 4 ZIP codes, in terms of number of active loans, in the greater Los Angeles area and leave to follow up work to train on the full dataset. The ZIP Codes are: 92677, with 109642 loans in Orange County; 93065, with 98673 loans in Simi Valley; 91709, with 95497 loans in Chino Hills; 92336 with 94794 loans in Fontana. We follow the same data filtering procedure as in [22], and use a subset with 50 of the features included there, we filter out loans if of the following features are not present: FICO score, Original balance, Initial interest rate, and Current State. Some other features we use include Original LTV, Unemployment Rate, Current Interest Rate, and National Mortgage Rate. We deal with missing data in other features with the missing indicator method, see, for example [17]. After filtering $117,523$ loans remain, each active for on average 45 months, for a total of around 5 million loan month transitions. Figure 11 shows the empirical transition probabilities across the dataset, and Figure 11 shows the transition counts on the sampled test set. Figure 12 shows the number of active loans, the prepayment rate, and the foreclosure rate over time on the dataset. For example, note the elevated foreclosure rates during and after the 2008 financial crisis. Table 12 shows the full set of features for the Set-Sequence model.
Figure 10: Empirical Transition Probabilities for the full top 4 ZIP codes on the combined train, validation and test dataset.
Figure 11: Top 4 ZIP transition counts on the test set January 2010 - December 2023.
Figure 12: Active Loans, Foreclosure Rate, and Prepayment Rate over time.
Table 12: Feature definitions and possible values. We require that current state, FICO score, Original balance, and Initial Interest Rate are available, and use a missing indicator when other features are unavailable. The feature vector including the missing indicators has dimension 50.
Feature Values
Current state Current, 30 Days Delinquent, 60 Days Delinquent, 90+ Days Delinquent,
Paid Off, Foreclosure, Real-Estate Owned
FICO score Continuous
Original balance Continuous
Initial interest rate Continuous
Original LTV Continuous
State unemployment rate Continuous
National mortgage rate Continuous
Current balance Continuous
Current interest rate Continuous
Scheduled monthly principal and interest Continuous
Scheduled principal Continuous
Days delinquent Continuous
Prime mortgage flag True, False
Convertible flag True, False
Pool insurance flag True, False
Insurance only flag True, False
Prepay penalty flag True, False
Negative amortization flag True, False
Time since origination $<1$ Year,  $<5$ Years,  $<10$ Years
Original term $<17$ Years
Number of 30-day delinquencies (past 12 months) 0–12
Number of 60-day delinquencies (past 12 months) 0–12
Number of 90+-day delinquencies (past 12 months) 0–12
Number of 30-day foreclosures (past 12 months) 0–12
Number of Current occurrences (past 12 months) 0–12
F.2 Model Fitting
We create an additional "non active" state that serves as a mask to ensure all loan sequences cover the full dataset. We sample loans in the following way in each batch: First, randomly pick a start time, then collect all active loans at that start time, and pick a subset N of them (set to 2500 in our experiments). We make this operation more efficient by first sorting the loans by origination date. Since the Set-Sequence model uses a context window of 50 time steps we sample using overlap between the train, validation, and test sets and only compute train gradients and performance metrics on the masked indices within the respective partition.
F.3 Additional Experiments Mortgage Risk Prediction
F.3.1 Transition Analysis
Figure 13 shows the one vs rest AUC (see Section 5.2 for the definition) conditioned on the initial state for the Set-Sequence model, averaged over 10 seeds of samples on the test set, where for each seed we sample 25 sequences of 2500 loans from the test set.
Figure 13: AUC Matrix for the Set-Sequence Model with 50 Features, averaged across 10 random seeds. In parenthesis we show the standard deviation of the AUC for that transition. We only show the AUC for transitions that happened at least 10 times for each seed.
F.3.2 Year-by-year results
In this subsection, we aim to understand the robustness of the Set-Sequence model across time for the mortgage risk task. To that end, we first describe our refitting approach with time.
Yearly Refitting
To emphasize recent data, we weight training windows with an exponential–decay schedule whose half-life is $\tau$ (e.g., 24 months). Define
$\alpha\;=\;\frac{\ln 2}{\tau},\qquad t_{\max}\;=\;\max_{i}t_{i}.$
Each window at time $t_{i}$ receives the (unnormalized) weight
$w_{i}\;=\;\exp\!\bigl{[}\alpha\,(t_{i}-t_{\max})\bigr{]}.$
Consequently, a window that is $\tau$ time units older than $t_{\max}$ has half the weight of the most recent window. During refitting we sample windows with probability proportional to $w_{i}$, biasing the training set toward more recent periods.
Figure 14 shows the improvement for the Set-Sequence model when using refitting method with $\tau=24$ months compared with not retraining and not weighting the retraining samples. We see a performance improvement with the weighted refitting.
Figure 14: Comparing three methods for data refitting: 1. not retraining 2. retraining with equal weight on all times, and 3. retraining with more weight on later times.
Results
We compare the Set-Sequence model, the Gated Selection model, and the models from  [22] in the setting where we do a base training for 15 epochs up until year 2002, and then for each epoch the train-set gets extended one year further. We save a checkpoint each year the model is extended.
In addition we up-weight the more recent train data, this is described in Section F.3.2. Figure 15 and Figure 16 shows the results, comparing the Set-Seq, Gated Selection (see Appendix B), NN, and Logistic model, all using the same retraining method, and where we evaluate on one year at a time.
Figure 15: Average AUC, Cross-Entropy, and 30dd-60dd AUC for the Set-Seq, NN, and Logistic model, all using the same retraining method, and where we evaluate on one year at a time.
Figure 16: The AUC per year for different transitions.
We see that the challenging years around the financial crisis affects the logistic model the most. However, the relative performance of the Set-Seq model and the NN model remain broadly the same, although it widens somewhat during 2009.
F.3.3 Interpretability
We show that the Gated Selection model is interpretable. To see the patterns learned in the gating matrix $G$ more clearly we first sort the rows/columns by the zip-code (white lines), then, within each ZIP-code, we sort by prime/subprime/unknown loan type (red lines). From Figure 17 we see that the selection matrix G learn to distinguish between these categories, which is aligned with the importance of these categorizations suggested in prior work. This means that the set-summary of a mortgage that, say is in ZIP code 1 and is subprime, will be a weighted sum, with the majority of the weight on other subprime mortgages in ZIP code 1. Another point to note here is that the gating matrix meaningfully change with economic cycles, so by considering the year 2018 instead of 2012 gives a different structure (highlighting for example different economic conditions evolving through time in the different ZIP-codes).
Figure 17: The selection matrix $G$, sorted by ZIP-code (white lines), and prime/subprime/unknown status (red lines).
Appendix G Additional Metric Details
Here we present formal definitions of metrics used in the paper. We are interested in considering these metrics for the class k corresponding to the absorbing (rare) class. Given access to true transition probabilities in our synthetic task framework, we establish test set evaluation metrics focused on the absorbing (rare) state. For this absorbing state $k$, we define the area under the receiver operating characteristic curve (AUC) metric as
$\operatorname{AUC}(k)=P(y_{1,k}>y_{2,k}|y_{1}=k,y_{2}\neq k)+\frac{1}{2}P(y_{1,k}=y_{2,k}|y_{1}=k,y_{2}\neq k),$
which characterizes the model’s discriminative capacity in identifying transitions to the absorbing state. By adding negligible noise to the predictions, all scores will be different and the second term will equal zero. Given a dataset of $n$ samples, $P$ of which belongs to class $K$ and $N$ that does not belong to class $K$, the AUC can be computed with
$AUC(k)=\frac{1}{NP}\sum_{i=1,j=1}^{n,n}1_{\hat{y}_{i,k}>\hat{y}_{j,k}}1_{y_{i}=k}1_{y_{j}\neq k}.$ (7)
This is more efficient than creating the ROC curve and computing the area underneath it. To quantify prediction accuracy for the absorbing state, we compute the correlation between predicted and true transition probabilities
$\operatorname{Corr}(p_{k},\hat{p}_{k})=\frac{\operatorname{Cov}(p_{k},\hat{p}_{k})}{\sqrt{\operatorname{Var}(p_{k})\cdot\operatorname{Var}(\hat{p}_{k})}}.$
The coefficient of determination $R^{2}$ measures the proportion of variance captured by predictions for transitions to the absorbing state
$R^{2}=1-\frac{\sum_{i=1}^{n}(p_{i,k}-\hat{p}_{i,k})^{2}}{\sum_{i=1}^{n}(p_{i,k}-\bar{p}_{k})^{2}},$
where $\bar{p}_{k}=\frac{1}{n}\sum_{i=1}^{n}p_{i,k}$. Our synthetic setup enables direct probability-based evaluation, rather than relying on predicted class labels. The Kullback-Leibler divergence measures the difference between the predicted and true probability distributions
$D_{KL}(p_{k}|\hat{p}_{k})=\frac{1}{n}\sum_{i=1}^{n}p_{i,k}\log\left(\frac{p_{i,k}}{\hat{p}{i,k}}\right).$ This metric is non-negative and equals zero if and only if $p_{i,k}=\hat{p}_{i,k}$ for all $i$. Unlike cross-entropy between the true labels and the predicted probabilities, the KL divergence between the predicted and true transition probabilities explicitly shows the deviation from a perfect model.